OpenStudy (anonymous):

I WILL GIVE MEDAL!!!!!!!! Given a soda can with a volume of 21 and a diameter of 6, what is the volume of a cone that fits perfectly inside the soda can?

OpenStudy (anonymous):

i will give medal

OpenStudy (anonymous):

7

OpenStudy (anonymous):

THANKS can you help me with one more

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

do i know why what

OpenStudy (anonymous):

\[V_{cyl}=\pi\cdot r^2 \cdot h\]\[V_{cone}=\frac{ 1 }{ 3 }\pi\cdot r^2\cdot h\]

OpenStudy (anonymous):

okay i get it

OpenStudy (anonymous):



OpenStudy (anonymous):

what's your other question?

OpenStudy (anonymous):

The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 375 mm3, what is the volume of the stack of pennies on the right in cubic millimeters?)

OpenStudy (anonymous):

what do you think it is?